Find values for m and b in the following system so that the solution to the system is (-3,4)
5x + 7y = b
mx + y = 22

Answer by rapaljer(4671)

If (-3,4) is a solution of this system of equations, then x=-3 and y=4. Just substitute these values of x and y into the two equations. How does that Glade commercial go in the the USA? "Plug it in, plug it in!!!"
5x + 7y = b
mx + y = 22

5(-3) + 7(4) = b
m(-3) + 4 = 22

The first equation solves for b:
-15 + 28 = b
13 = b

The second equation solves for m:
-3m + 4 = 22
-3m = 18
m=-6

R^2 at SCC